OpenStudy (anonymous):

5x + 3y = -2. Find the rate of change.

OpenStudy (anonymous):

rate of change is just the slope of the line.

OpenStudy (anonymous):

you can solve for y then the number before the x is your slope

OpenStudy (anonymous):

5x+3y=-2 Since 5x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 5x from both sides. 3y=-5x-2 Divide each term in the equation by 3. (3y)/(3)=-(5x)/(3)-(2)/(3) Cancel the common factor of 3 in (3y)/(3). (<X>3<x>y)/(<X>3<x>)=-(5x)/(3)-(2)/(3) Remove the common factors that were cancelled out. y=-(5x)/(3)-(2)/(3) To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept. y=mx+b Using the y=mx+b formula, m=-(5)/(3) and b=-(2)/(3). m=-(5)/(3), b=-(2)/(3)

OpenStudy (anonymous):

now that you have the slope and y-intercept do you know how to proceed and get your answer?